# PEneo: Unifying Line Extraction, Line Grouping, and Entity Linking for End-to-end Document Pair Extraction

### v1

@@ 5.3 Baseline Settings
LiLT[InfoXLM] [23] | 66.50 | 43.98 | 63.85 | 62.60 | 60.57 | 55.13 | 52.96 |
PEneo-LiLT[InfoXLM] | 80.51 (+14.01) | 54.59 (+10.61) | 71.43 (+7.58) | 77.49 (+14.89) | 73.62 (+13.05) | 70.11 (+14.98) | 71.43 (+18.47) |
LayoutXLM [27] | 64.11 | 40.21 | 66.75 | 67.98 | 63.04 | 58.77 | 59.79 |
PEneo-LayoutXLM | 80.41 (+16.30) | 52.81 (+12.60) | 74.56 (+7.81) | 78.11 (+10.13) | 75.17 (+12.13) | 74.06 (+15.29) | 70.81 (+11.02) |
5.4 Comparison with Existing Methods
Results are shown in Table 2-4. Previous pipelines underperform on the pair extraction task. Donut utilizes the visual modality only, which may be insufficient for analyzing the diverse and complex key-value relationships in the document. We argue that Donut performs better in scenarios with a small number of output tokens, such as receipt understanding in CORD [20]. Its prediction ability is not yet fully developed for documents with a large number of texts. The SER+RE pipelines suffer from performance drop, mainly due to the error accumulation between modules and the improper text order, which will be discussed in the subsequent sections. PEneo, on the other hand, substantially improves the performance of each backbone. On RFUND-EN, the F1 score of the entire pipeline is boost by 19.89% for LiLT[EN-R], 22.11% for LiLT[InfoXLM], 21.27% for LayoutXLM, 22.91% for LayoutLMv2, and 21.61% for LayoutLMv3. Most of these backbones outperform the strong baseline GeoLayoutLM which contains task-specific pre-trained RE modules, although there exist huge gaps between them under the previous SER+RE setting. For the other language subset of RFUND, PEneo still offers substantial performance improvements. On the SIBR dataset, the new pipeline has demonstrated a score improvement ranging from 9.01% to 11.78%, confirming its ability in bilingual settings. These outcomes underscore the effectiveness and versatility of PEneo, as it consistently achieves performance gains across multiple language scenarios and diverse backbone configurations.
5.5 Analysis of Module Collaboration
@@
We observe that LiLT and LayoutLMv3 underperform in several cases, which seems to be contradictory to the results reported in previous literature [23, 10] at first glance. In fact, these two methods utilize entity-level boxes for layout modeling, while the conventional settings ([26, 28, 27, 19]) use word-level boxes. In our experiment, all the models take the line-level coordinates as input, which affects their SER ability to some extent.
To further illustrate the performance drop phenomenon in the SER+RE pipeline, we conducted experiments on RFUND-EN with LiLT[InfoXLM] using different SER results. As shown in Table 5, a performance gap of 19 points in the SER module results in a 15-point decrease in pair extraction performance (#1 and #2). As the SER score is further reduced by 3 points, the performance of the pair extraction reduces by 3 points as well (#2 and # 3, #4). Results show that the SER+RE approach exhibits a serious accumulation of errors. The accuracy of the SER module greatly affects the performance of the whole pipeline.
We then explore the impact of different types of SER errors on the whole pipeline. Commonly seen errors in the SER step include: (1) entity false negative, where keys and values are categorized as background elements; (2) entity false positive, where background elements are categorized as keys or values; (3) entity category error, where keys/values are identified as values/keys; (4) entity fragmentation, where an entity is recognized as multiple parts belong to different categories, as shown in Figure 3(a). We added the above disturbances to the SER ground truths with different probabilities and tested the performance variation of pair extraction under the setting of a fixed RE module. Results are shown in Figure 5. Entity fragmentation has the largest effect since inaccurate entity spans inherently lead to mistakes. False positives have the smallest influence as the RE module could filter out some misclassifications given accurate spans. False negatives directly cause missing links, while category errors interfere with the RE module’s reasoning. Overall, precise entity span detection proved critical for the SER+RE pipeline. However, it is highly dependent on the correct order of input text and large granularity of coordinate information [16], which is often difficult to achieve in practice, making the model underperform.
@@
6 Conclusion and Future Work

### v2

@@ 5.3. Baseline Settings
LiLT[InfoXLM]BASE (Wang et al., 2022) | 66.50 | 43.98 | 63.85 | 62.60 | 60.57 | 55.13 | 52.96 |
PEneo-LiLT[InfoXLM]BASE | 80.51 (+14.01) | 54.59 (+10.61) | 71.43 (+7.58) | 77.49 (+14.89) | 73.62 (+13.05) | 70.11 (+14.98) | 71.43 (+18.47) |
LayoutXLMBASE (Xu et al., 2021a) | 64.11 | 40.21 | 66.75 | 67.98 | 63.04 | 58.77 | 59.79 |
PEneo-LayoutXLMBASE | 80.41 (+16.30) | 52.81 (+12.60) | 74.56 (+7.81) | 78.11 (+10.13) | 75.17 (+12.13) | 74.06 (+15.29) | 70.81 (+11.02) |
5.4. Comparison with Existing Methods
Results are shown in Table 2-4. Previous pipelines underperform on the pair extraction task. FUDGE and Donut utilize the visual modality only, which may be insufficient for analyzing the diverse and complex key-value relationships in the document. We argue that Donut performs better in scenarios with a small number of output tokens, such as receipt understanding in CORD (Park et al., 2019). Its prediction ability is not yet fully developed for documents with a large number of texts. TPP employs a unified token-path-prediction pipeline to tackle the RE task, without a dedicated design for error suppression. In this schema, even a minor error in the linking prediction matrix could result in a completely incorrect outcome. For GPT-4V, integrating the OCR results into the prompt can significantly improve its performance, but it still falls short of numerous supervised approaches. Upon analyzing its output, we argue that its underperformance can be attributed primarily to the LLM hallucination, which leads to numerous redundant or inaccurate predictions. The SER+RE pipelines suffer from performance drop, mainly due to the error accumulation between modules and the improper text order, which will be discussed in the subsequent sections. PEneo, on the other hand, substantially improves the performance of each backbone. On RFUND-EN, the F1 score of the entire pipeline is boost by 19.89% for LiLT[EN-R]BASE, 22.11% for LiLT[InfoXLM]BASE, 21.27% for LayoutXLMBASE, 22.91% for LayoutLMv2BASE, and 21.61% for LayoutLMv3BASE. Most of these backbones outperform the strong baseline GeoLayoutLM which contains task-specific pre-trained RE modules, although there exist huge gaps between them under the previous SER+RE setting. For the other language subset of RFUND, PEneo still offers substantial performance improvements. On the SIBR dataset, the new pipeline has demonstrated a score improvement ranging from 9.01% to 11.78%, confirming its ability in bilingual settings. These outcomes underscore the effectiveness and versatility of PEneo, as it consistently achieves performance gains across multiple language scenarios and diverse backbone configurations.
5.5. Analysis of Module Collaboration
@@
We observe that LiLT and LayoutLMv3 underperform in several cases, which seems to be contradictory to the results reported in previous literature (Wang et al., 2022; Huang et al., 2022) at first glance. In fact, these two methods utilize entity-level boxes for layout modeling, while the conventional settings (Xu et al., 2020, 2021b, 2021a; Luo et al., 2023) use word-level boxes. In our experiment, all the models take the line-level coordinates as input, which affects their SER ability to some extent.
To further illustrate the performance drop phenomenon in the SER+RE pipeline, we conducted experiments on RFUND-EN with LiLT[InfoXLM]BASE using different SER results. As shown in Table 5, a performance gap of 19 points in the SER module results in a 15-point decrease in pair extraction performance (#1 and #2). As the SER score is further reduced by 3 points, the performance of the pair extraction reduces by 3 points as well (#2 and # 3, #4). Results show that the SER+RE approach exhibits a serious accumulation of errors. The accuracy of the SER module greatly affects the performance of the whole pipeline.
We then explore the impact of different types of SER errors on the whole pipeline. Commonly seen errors in the SER step include: (1) entity false negative, where keys and values are categorized as background elements; (2) entity false positive, where background elements are categorized as keys or values; (3) entity category error, where keys/values are identified as values/keys; (4) entity fragmentation, where an entity is recognized as multiple parts belong to different categories, as shown in Figure 4(a). We added the above disturbances to the SER ground truths with different probabilities and tested the performance variation of pair extraction under the setting of a fixed RE module. Results are shown in Figure 5. Entity fragmentation has the largest effect since inaccurate entity spans inherently lead to mistakes. False positives have the smallest influence as the RE module could filter out some misclassifications given accurate spans. False negatives directly cause missing links, while category errors interfere with the RE module’s reasoning. Overall, precise entity span detection proved critical for the SER+RE pipeline. However, it is highly dependent on the correct order of input text and large granularity of coordinate information (Li et al., 2021b), which is often difficult to achieve in practice, making the model underperform.
@@
6. Conclusion and Future Work
@@
Acknowledgements.
This research is supported in part by National Natural Science Foundation of China (Grant No.: 62441604, 61936003).References
- (1)
- Achiam et al. (2023) Josh Achiam, Steven Adler, Sandhini Agarwal, Lama Ahmad, Ilge Akkaya, Florencia Leoni Aleman, Diogo Almeida, Janko Altenschmidt, Sam Altman, Shyamal Anadkat, et al. 2023. Gpt-4 technical report. arXiv preprint arXiv:2303.08774 (2023).
- Appalaraju et al. (2021) Srikar Appalaraju, Bhavan Jasani, Bhargava Urala Kota, Yusheng Xie, and R Manmatha. 2021. DocFormer: End-to-End Transformer for Document Understanding. In 2021 IEEE/CVF International Conference on Computer Vision (ICCV). IEEE, 973–983.
- Cao et al. (2022) Haoyu Cao, Xin Li, Jiefeng Ma, Deqiang Jiang, Antai Guo, Yiqing Hu, Hao Liu, Yinsong Liu, and Bo Ren. 2022. Query-Driven Generative Network for Document Information Extraction in the Wild. In Proceedings of the 30th ACM International Conference on Multimedia. 4261–4271.
- Dai et al. (2019) Dai Dai, Xinyan Xiao, Yajuan Lyu, Shan Dou, Qiaoqiao She, and Haifeng Wang. 2019. Joint Extraction of Entities and Overlapping Relations Using Position-Attentive Sequence Labeling. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 33. 6300–6308.
- Davis et al. (2022) Brian Davis, Bryan Morse, Brian Price, Chris Tensmeyer, Curtis Wigington, and Vlad Morariu. 2022. End-to-End Document Recognition and Understanding with Dessurt. In European Conference on Computer Vision. Springer, 280–296.
- Davis et al. (2021) Brian Davis, Bryan Morse, Brian Price, Chris Tensmeyer, and Curtis Wiginton. 2021. Visual FUDGE: Form Understanding via Dynamic Graph Editing. In Document Analysis and Recognition–ICDAR 2021: 16th International Conference, Lausanne, Switzerland, September 5–10, 2021, Proceedings, Part I 16. Springer, 416–431.
- Gu et al. (2021) Jiuxiang Gu, Jason Kuen, Vlad I Morariu, Handong Zhao, Rajiv Jain, Nikolaos Barmpalios, Ani Nenkova, and Tong Sun. 2021. UniDoc: Unified Pretraining Framework for Document Understanding. Advances in Neural Information Processing Systems 34 (2021), 39–50.
- Gu et al. (2022) Zhangxuan Gu, Changhua Meng, Ke Wang, Jun Lan, Weiqiang Wang, Ming Gu, and Liqing Zhang. 2022. XYLayoutLM: Towards Layout-aware Multimodal Networks for Visually-rich Document Understanding. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 4583–4592.
- Guillaume Jaume (2019) Jean-Philippe Thiran Guillaume Jaume, Hazim Kemal Ekenel. 2019. FUNSD: A Dataset for Form Understanding in Noisy Scanned Documents. In ICDAR-OST.
- Hong et al. (2022) Teakgyu Hong, Donghyun Kim, Mingi Ji, Wonseok Hwang, Daehyun Nam, and Sungrae Park. 2022. BROS: A Pre-trained Language Model Focusing on Text and Layout for Better Key Information Extraction from Documents. In Proceedings of the AAAI Conference on Artificial Intelligence, Vol. 36. 10767–10775.
- Hu et al. (2023) Kai Hu, Zhuoyuan Wu, Zhuoyao Zhong, Weihong Lin, Lei Sun, and Qiang Huo. 2023. A Question-Answering Approach to Key Value Pair Extraction from Form-Like Document Images. Proceedings of the AAAI Conference on Artificial Intelligence 37, 11 (Jun. 2023), 12899–12906.
- Huang et al. (2022) Yupan Huang, Tengchao Lv, Lei Cui, Yutong Lu, and Furu Wei. 2022. LayoutLMv3: Pre-training for Document AI with Unified Text and Image Masking. In Proceedings of the 30th ACM International Conference on Multimedia. 4083–4091.
- Huang et al. (2019) Zheng Huang, Kai Chen, Jianhua He, Xiang Bai, Dimosthenis Karatzas, Shijian Lu, and CV Jawahar. 2019. ICDAR2019 Competition on Scanned Receipt OCR and Information Extraction. In 2019 International Conference on Document Analysis and Recognition (ICDAR). IEEE, 1516–1520.
- Hwang et al. (2021) Wonseok Hwang, Jinyeong Yim, Seunghyun Park, Sohee Yang, and Minjoon Seo. 2021. Spatial Dependency Parsing for Semi-Structured Document Information Extraction. In Findings of the Association for Computational Linguistics: ACL-IJCNLP 2021. 330–343.
- Kim et al. (2022) Geewook Kim, Teakgyu Hong, Moonbin Yim, JeongYeon Nam, Jinyoung Park, Jinyeong Yim, Wonseok Hwang, Sangdoo Yun, Dongyoon Han, and Seunghyun Park. 2022. OCR-free Document Understanding Transformer. In European Conference on Computer Vision. Springer, 498–517.
- Kuang et al. (2023) Jianfeng Kuang, Wei Hua, Dingkang Liang, Mingkun Yang, Deqiang Jiang, Bo Ren, and Xiang Bai. 2023. Visual information extraction in the wild: practical dataset and end-to-end solution. In International Conference on Document Analysis and Recognition. Springer, 36–53.
- Li et al. (2021a) Peizhao Li, Jiuxiang Gu, Jason Kuen, Vlad I Morariu, Handong Zhao, Rajiv Jain, Varun Manjunatha, and Hongfu Liu. 2021a. SelfDoc: Self-Supervised Document Representation Learning. In 2021 IEEE/CVF Conference on Computer Vision and Pattern Recognition (CVPR). IEEE, 5648–5656.
- Li et al. (2021b) Yulin Li, Yuxi Qian, Yuechen Yu, Xiameng Qin, Chengquan Zhang, Yan Liu, Kun Yao, Junyu Han, Jingtuo Liu, and Errui Ding. 2021b. StrucTexT: Structured Text Understanding with Multi-modal Transformers. In Proceedings of the 29th ACM International Conference on Multimedia. 1912–1920.
- Liao et al. (2023) Haofu Liao, Aruni RoyChowdhury, Weijian Li, Ankan Bansal, Yuting Zhang, Zhuowen Tu, Ravi Kumar Satzoda, R Manmatha, and Vijay Mahadevan. 2023. DocTr: Document transformer for structured information extraction in documents. In Proceedings of the IEEE/CVF International Conference on Computer Vision. 19584–19594.
- Loshchilov and Hutter (2019) Ilya Loshchilov and Frank Hutter. 2019. Decoupled Weight Decay Regularization. In Proceedings of the 7th International Conference on Learning Representations (ICLR).
- Luo et al. (2023) Chuwei Luo, Changxu Cheng, Qi Zheng, and Cong Yao. 2023. GeoLayoutLM: Geometric Pre-training for Visual Information Extraction. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 7092–7101.
- Park et al. (2019) Seunghyun Park, Seung Shin, Bado Lee, Junyeop Lee, Jaeheung Surh, Minjoon Seo, and Hwalsuk Lee. 2019. CORD: A Consolidated Receipt Dataset for Post-OCR Parsing. In Workshop on Document Intelligence at NeurIPS 2019.
- Peng et al. (2022) Qiming Peng, Yinxu Pan, Wenjin Wang, Bin Luo, Zhenyu Zhang, Zhengjie Huang, Yuhui Cao, Weichong Yin, Yongfeng Chen, Yin Zhang, et al. 2022. ERNIE-Layout: Layout Knowledge Enhanced Pre-training for Visually-rich Document Understanding. In Findings of the Association for Computational Linguistics: EMNLP 2022. 3744–3756.
- Ramshaw and Marcus (1999) Lance A Ramshaw and Mitchell P Marcus. 1999. Text chunking using transformation-based learning. In Natural language processing using very large corpora. Springer, 157–176.
- Seki et al. (2007) Minenobu Seki, Masakazu Fujio, Takeshi Nagasaki, Hiroshi Shinjo, and Katsumi Marukawa. 2007. Information Management System Using Structure Analysis of Paper/Electronic Documents and Its Applications. In Ninth International Conference on Document Analysis and Recognition (ICDAR 2007), Vol. 2. 689–693. https://doi.org/10.1109/ICDAR.2007.4377003
- Shi et al. (2023) Yongxin Shi, Dezhi Peng, Wenhui Liao, Zening Lin, Xinhong Chen, Chongyu Liu, Yuyi Zhang, and Lianwen Jin. 2023. Exploring ocr capabilities of gpt-4v (ision): A quantitative and in-depth evaluation. arXiv preprint arXiv:2310.16809 (2023).
- Soares et al. (2019) Livio Baldini Soares, Nicholas Fitzgerald, Jeffrey Ling, and Tom Kwiatkowski. 2019. Matching the Blanks: Distributional Similarity for Relation Learning. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics. 2895–2905.
- Wang et al. (2019) Haoyu Wang, Ming Tan, Mo Yu, Shiyu Chang, Dakuo Wang, Kun Xu, Xiaoxiao Guo, and Saloni Potdar. 2019. Extracting Multiple-Relations in One-Pass with Pre-Trained Transformers. In Proceedings of the 57th Annual Meeting of the Association for Computational Linguistics. 1371–1377.
- Wang et al. (2022) Jiapeng Wang, Lianwen Jin, and Kai Ding. 2022. LiLT: A Simple yet Effective Language-Independent Layout Transformer for Structured Document Understanding. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 7747–7757.
- Wang et al. (2020) Yucheng Wang, Bowen Yu, Yueyang Zhang, Tingwen Liu, Hongsong Zhu, and Limin Sun. 2020. TPLinker: Single-stage Joint Extraction of Entities and Relations Through Token Pair Linking. In Proceedings of the 28th International Conference on Computational Linguistics. 1572–1582.
- Watanabe et al. (1995) T. Watanabe, Qin Luo, and N. Sugie. 1995. Layout Recognition of Multi-Kinds of Table-Form Documents. IEEE Transactions on Pattern Analysis and Machine Intelligence 17, 4 (1995), 432–445. https://doi.org/10.1109/34.385976
- Wei et al. (2020) Zhepei Wei, Jianlin Su, Yue Wang, Yuan Tian, and Yi Chang. 2020. A Novel Cascade Binary Tagging Framework for Relational Triple Extraction. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics. 1476–1488.
- Xu et al. (2020) Yiheng Xu, Minghao Li, Lei Cui, Shaohan Huang, Furu Wei, and Ming Zhou. 2020. LayoutLM: Pre-training of Text and Layout for Document Image Understanding. In Proceedings of the 26th ACM SIGKDD International Conference on Knowledge Discovery & Data Mining. 1192–1200.
- Xu et al. (2021a) Yiheng Xu, Tengchao Lv, Lei Cui, Guoxin Wang, Yijuan Lu, Dinei Florencio, Cha Zhang, and Furu Wei. 2021a. LayoutXLM: Multimodal Pre-training for Multilingual Visually-rich Document Understanding. arXiv:2104.08836 [cs.CL]
- Xu et al. (2022) Yiheng Xu, Tengchao Lv, Lei Cui, Guoxin Wang, Yijuan Lu, Dinei Florencio, Cha Zhang, and Furu Wei. 2022. XFUND: A Benchmark Dataset for Multilingual Visually Rich Form Understanding. In Findings of the Association for Computational Linguistics: ACL 2022. 3214–3224.
- Xu et al. (2021b) Yang Xu, Yiheng Xu, Tengchao Lv, Lei Cui, Furu Wei, Guoxin Wang, Yijuan Lu, Dinei Florencio, Cha Zhang, Wanxiang Che, et al. 2021b. LayoutLMv2: Multi-modal Pre-training for Visually-rich Document Understanding. In Proceedings of the 59th Annual Meeting of the Association for Computational Linguistics and the 11th International Joint Conference on Natural Language Processing (Volume 1: Long Papers). 2579–2591.
- Yang et al. (2023) Zhibo Yang, Rujiao Long, Pengfei Wang, Sibo Song, Humen Zhong, Wenqing Cheng, Xiang Bai, and Cong Yao. 2023. Modeling Entities as Semantic Points for Visual Information Extraction in the Wild. In Proceedings of the IEEE/CVF Conference on Computer Vision and Pattern Recognition. 15358–15367.
- Ye et al. (2022) Deming Ye, Yankai Lin, Peng Li, and Maosong Sun. 2022. Packed Levitated Marker for Entity and Relation Extraction. In Proceedings of the 60th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). 4904–4917.
- Zhang et al. (2023) Chong Zhang, Ya Guo, Yi Tu, Huan Chen, Jinyang Tang, Huijia Zhu, Qi Zhang, and Tao Gui. 2023. Reading Order Matters: Information Extraction from Visually-rich Documents by Token Path Prediction. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing. 13716–13730.
- Zheng et al. (2017) Suncong Zheng, Feng Wang, Hongyun Bao, Yuexing Hao, Peng Zhou, and Bo Xu. 2017. Joint Extraction of Entities and Relations Based on a Novel Tagging Scheme. In Proceedings of the 55th Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers). Association for Computational Linguistics, Vancouver, Canada, 1227–1236. https://doi.org/10.18653/v1/P17-1113
- Zhong and Chen (2021) Zexuan Zhong and Danqi Chen. 2021. A Frustratingly Easy Approach for Entity and Relation Extraction. In Proceedings of the 2021 Conference of the North American Chapter of the Association for Computational Linguistics: Human Language Technologies. 50–61.